### v4

@@ Abstract
Wide-angle lenses are commonly used in perception tasks requiring a large field of view. Unfortunately, these lenses produce significant distortions, making conventional models that ignore the distortion effects unable to adapt to wide-angle images. In this paper, we present a novel transformer-based model that automatically adapts to the distortion produced by wide-angle lenses. Our proposed image encoder architecture, dubbed , leverages the physical characteristics of such lenses analytically defined by the radial distortion profile. In contrast to conventional transformer-based architectures, comprises a radial patch partitioning, a distortion-based sampling technique for creating token embeddings, and an angular position encoding for radial patch merging. Compared to other baselines, achieves the best results on different datasets with significant gains when trained on bounded levels of distortions (very low, low, medium, and high) and tested on all, including out-of-distribution distortions. While the base architecture requires knowledge of the radial distortion profile, we show it can be combined with a self-calibration network that estimates such a profile from the input image itself, resulting in a completely uncalibrated pipeline. Finally, we also present DarSwin-Unet, which extends to an encoder-decoder architecture suitable for pixel-level tasks. We demonstrate its performance on depth estimation and show through extensive experiments that DarSwin-Unet can perform zero-shot adaptation to unseen distortions of different wide-angle lenses.
@@ 1 Introduction
Figure 1: Illustration of (cartesian) Swin [5] (left) and our (radial) (right) given a wide-angle image (middle). While Swin [5] computes attention on the predefined windows over square image patches (bottom left orange region), performs radial transformations using distortion-aware radial patches and computes the attention on windows defined over radial patches (shown in orange region bottom right), which enables greater generalization capabilities across different lenses.
@@
Unfortunately, such wide-angle lenses create significant distortions in the image since the perspective projection model no longer applies: straight lines appear curved, and the appearance of the same object changes as a function of its position on the image plane. This distortion breaks the translation equivariance assumption implicit in convolution neural networks (CNNs) and, therefore, limits their applicability. This problem is further exacerbated by the diversity in lens distortion profiles: as we will demonstrate, a network naively trained for a specific lens tends to overfit to that specific distortion and does not generalize well when tested on another lens. Just as methods are needed to address the “domain gap” [6] from dataset bias [7, 8], we believe we must also bridge the “distortion gap” to truly make wide-angle imaging applicable.
@@
One popular strategy to bridge the distortion gap is to cancel the effect of distortion on the image plane by warping the input image back to a perspective projection model according to calibrated lens parameters. Conventional approaches can then be trained and tested on the resulting “un-distorted” images. A wide array of such methods, ranging from classical [9, 10, 11, 12, 13] to deep learning [14, 15], have been proposed. Unfortunately, warping a very wide angle image to a perspective projection tends to create severely stretched images and restricts the maximum field of view since, in the limit, a point at 90°90°90  ° azimuth projects at infinity. Reducing the maximum field of view defeats the purpose of using a wide-angle lens in the first place. Other projections are also possible (e.g., cylindrical [16], or piecewise linear [4]), but these also tend to create unwanted distortions or suffer from resolution loss.
@@
Recently, methods that break free from the “undistort first” strategy aim to reason directly about the images without undistorting them. For example, methods like [17, 18] use deformable convolutions [19, 20] to adapt convolution kernels to the lens distortion. However, the high computational cost of deformable CNNs constrains the kernel adaptation to a few layers inside the network. Other approaches like spherical CNNs [21] or gauge equivariant CNN [22] can adapt to different manifolds, but their applicability for lens distortion has not been demonstrated. Finally, vision transformers [23] and their more recent variants [24, 5, 25] could also better bridge the “distortion gap” since they do not assume any prior structure other than permutation equivariance, but their cartesian partitioning of the image plane does not take lens geometry into account (see fig. 1).
@@
In this paper, we present , a transformer-based image encoder architecture that adapts its structure to the lens distortion profile, which is assumed to be known (calibrated lens). Our method, inspired by the recent Swin transformer architecture [5], leverages a distortion-aware sampling scheme for creating token embeddings, employs polar patch partitioning and merging strategies, and relies on angular relative positional encoding. This explicitly embeds knowledge of the lens distortion in the architecture and makes it much more robust to the “distortion gap” created by training and testing on different lenses. We show, through extensive classification experiments, that can perform zero-shot adaptation (without pretraining) across different lenses compared to the baselines Swin [5] trained on distorted and undistorted images and DAT [25] trained on distorted images.
While the encoder architecture was originally introduced in [26], this paper significantly extends the original contribution in three main ways. First, we show that, by combining a DarSwin encoder with another network that estimates the lens distortion from the image, we can remove the need for calibrating the lens profile, resulting in an entirely uncalibrated pipeline. Second, we extend the DarSwin encoder to a full encoder-decoder distortion-aware architecture, DarSwin-Unet, suitable for per-pixel prediction tasks. Third, we present experiments on the depth estimation task, which show that DarSwin-Unet is much more robust to changes in lens distortions at test time than all of the compared baselines, including Swin-Unet [27] and Swin-UPerNet [5, 28] trained on both distorted and undistorted images, and DAT-UPerNet [25, 28] trained on distorted images.
@@ 2 Related work
Panoramic images span the full 360°360°360  ° field of view and are most commonly projected onto a plane using an equirectangular projection. This projection introduces significant distortions, especially near the poles. Many works have explicitly designed approaches to deal with equirectangular distortion, to perform different downstream tasks like depth estimation [29], saliency detection [30], segmentation [31], layout estimation [32], and object detection [33] to name a few. However, these are specifically designed for equirectangular distortion and do not generalize to wide-angle lenses.
@@
Establishing image and object recognition tasks from wide-angle cameras directly is relatively recent [34, 35, 16, 36, 4] due to the presence of distortions in the image. In computer vision, the application of wide-angle images ranges from visual perception [37] to autonomous vehicle cameras [4, 38, 39]. In this respect, the initial studies mainly focused on correcting the distortion of the image [15, 40, 14, 11, 41, 42, 43, 44, 45, 46]. Many methods like [42, 47, 43] use a distortion distribution map to correct distortion. Distortion correction can lead to multiple types of sampling artifacts, as explained in [4], leading to a performance loss. In contrast, relies on distortion parameters and reasons directly on wide-angle images.
@@
CNNs [48, 49, 50] are particularly well-suited for perspective images due to their implicit bias and translational equivariance [51]. Methods like [52, 35, 34] try to adapt CNNs on fisheye images for tasks such as object detection. However, the distortion caused by wide-angle images breaks this symmetry, which reduces the generalization performance of CNNs. Methods like [53, 54, 55, 56] use self-supervised learning combined with techniques like distillation or multi-task learning to have a better understanding of distortion. Deformable convolutions [19] (later extended in [20]) learn a deformation to be applied to convolution kernels, which can provide greater flexibility at the cost of significant additional computation. Closer to our work, [18, 17, 3, 57] use such deformable CNNs to understand the distortion in a fisheye image. In contrast, leverages attention-based mechanisms rather than convolution.
@@
Vision transformers (ViT) [23] use self-attention mechanisms [58] computed on image patches rather than performing convolutions. Unlike CNNs, a ViT does not have a fixed geometric structure in its architecture: any extra structure is given via positional encoding. More recently, the Swin transformer architecture [5] proposes a multi-scale strategy of window-based attention. Later, deformable attention transformer (DAT) [25], adapts the concept of deformable CNNs [19, 20] to increase adaptability. [59] proposes a distortion-aware architecture using a transformer network, but the network is limited to a fixed equirectangular distortion. In contrast, adapts its structure to the lens distortion, which may vary across lenses.
Hybrid-network based approaches
Some methods leverage properties from both self-attention and convolutions and build hybrid networks. Methods like [60] use hybrid networks and try to leverage the geometric property of fisheye images (i.e., the orthogonal placement of objects) and propose a new representation of fisheye road scenes, invariant to the camera viewing direction. Shi et al. [61] leverages the radial nature of distortion by including polar cross attention for inpainting, but unlike , they do not use the lens information in their network. Similar to our method, [56, 54, 62, 63] propose a camera-aware depth estimation network to handle the severe distortion of fisheye cameras: [54] encode the camera intrinsic parameters as a tensor; and [63] propose a self-supervised depth estimation method which relies on the lens distortion parameter for forward and back-projection functions. Both these methods use distortion parameters as a part of the input or training process, but they rely on convolutions whose generalization capabilities are limited due to the translational invariance assumption being broken in wide-angle images. Indeed, the network weights each pixel and its corresponding lens distortion prior equally regardless of the severity of distortion, which affects the ability to generalize to a variety of lens distortion. Hence, [54] shows generalization to lens distortion closer to training distortion, and [63] does not show any generalization results. In contrast, our approach uses the lens parameters to adapt the structure of the input explicitly, weighting each part of the image according to the severity of lens distortion. Hence, we show generalization to a vast range of distortion parameters.
@@ 3 Image formation
Figure 2: Overview of our distortion aware transformer unet architecture, DarSwin-Unet which. It employs hierarchical layers of ’s transformer blocks [26] (top row) and is inspired by unet architecture from Swin-Unet [27]. To make it adapt to lens distortion, the patch partition, linear embedding, and patch merging layers and patch expanding layer all take the lens projection curve $\mathcal{P}$ (c.f. sec. 3) as input similar to . The $k$-NN layer is used to project the feature map from polar ($N_{r}\times N_{\varphi})$ to cartesian space $H\times W$.
@@ 4.1 Overview
fig. 2 shows an overview of our distortion-aware transformer model. This model extends the encoder architecture (), proposed in [26] for classification tasks, to the new DarSwin-Unet architecture that scales up to pixel-level tasks. The encoder () accepts as input a single image and its distortion parameters in the form of a lens projection curve $\mathcal{P}(\theta)$ (see eq. 2). The image is first segmented into patches according to a polar partitioning module (sec. 4.2). Then, a linear embedding is computed from sampled points (sec. 4.3), reshaped into a radial-azimuth projection and fed to the first Swin transformer block. The attention mechanism employs an angular relative positional encoding scheme guided by the lens curve (sec. 4.5). This is followed by three blocks performing patch merging (sec. 4.6) and additional Swin transformer blocks. The decoder employs similar transformer blocks along with azimuth patch expanding layer for upsampling (sec. 4.7). A $k$-NN layer (sec. 4.8) then projects the feature map from polar to cartesian space and is followed by a final linear projection layer. More details on each of these steps are provided below.
@@
Figure 3: Illustration of the difference between (a) Swin [5], which uses Cartesian image patches (shown in grey) and windows (shown in white borders) by making divisions along image axes, Swin performs attention on a Cartesian window and merges 2 $\times$ 2 neighbourhood patches to build a hierarchical structure (shown in the green and orange shaded region) (b) our proposed uses polar image patches by making divisions along radius and azimuth. In our case, the images patches (shown in grey) and windows (shown in white borders) are defined along azimuth and merged along azimuth as (shown in the green and orange shaded region)
@@ 4.2 Polar partition
Figure 4: Example of distortion aware polar patch partition. Here, $N_{r}=4$ and $N_{\varphi}=4$ partitions along radius and azimuth, respectively, are used to illustrate (right, only the top-right quadrant of the image is shown). While the azimuth partitioning is performed in an equiangular fashion, the radial dimension takes the lens distortion curve (left) into account. The field of view along the incident angle $\theta$ is split into $N_{r}$ equal parts (left), and corresponding radial are obtained from the distortion curve. Hence, for different lenses, we can have different radii depending on the distortion parameters.
The first step of our proposed architecture is to partition the image domain into patches. As opposed to Swin, which performs the split in cartesian coordinates (fig. 3a), employs a polar patch partitioning strategy (fig. 3b). After centering a polar coordinate system on the image center (assumed to be known), we first split according to azimuth angle $\varphi$ in $N_{\varphi}$ equiangular regions. For the radial dimension, we split the image into $N_{r}$ radial regions such that the splits are equiangular in $\theta$ and obtain the corresponding image-space radii using the lens projection function $\mathcal{P}(\theta)$ (see eq. 2). The total number of patches is therefore $N_{\varphi}\times N_{r}$. In our experiments, we set $N_{r}=16$ and $N_{\varphi}=64$ for an input image of size $64\times 64$ pixels.
@@ 4.3 Linear embedding
The resulting image patches created by this approach have an unequal amount of pixels. Therefore, we rely on a distortion aware sampling strategy to obtain the same number of points for each patch. To sample from the images, we define the number of sampling points along the radius and azimuth, as shown in fig. 5, and adapt the pattern according to each partition. In our experiments, we set 10 sampling points along the radius and azimuth for each patch. Points are sampled in an equiangular fashion along the azimuth direction. For the radial dimension, we sample according to the same pattern as the polar partitioning (sec. 4.2); that is, we split in an equiangular fashion according to $\theta$ and obtain the corresponding image-space radii using $\mathcal{P}(\theta)$ (eq. 2). The input image is sampled using bilinear interpolation, and samples are arranged in a polar format, as illustrated in fig. 4. The resulting sample values are then fed into a linear embedding layer to produce token embeddings as input for the subsequent transformer block.
Figure 5: Example of sampling strategy on patch partitions. For illustration, 16 patches ($N_{r}=2$ and $N_{\varphi}=8$) are used, with window size $(1,2)$, i.e. ($M_{r}=1,M_{\varphi}=2$). RGB values are then sampled bilinearly from the image (6 blue dots per patch) and re-arranged in polar coordinates.
@@ 4.4 Window-based self attention
The Swin transformer architecture [5] uses window-based self-attention, where attention is computed on non-overlapping windows of $M\times M$ patches. Here, we maintain that strategy, but the polar nature of our patches allows for an additional design choice: the number of patches along azimuth $M_{\varphi}$ and radius $M_{r}$ can be different. As shown in tab. 1, we define two variants: $M$ patches along both azimuth and radius, $M_{r}=M_{\varphi}=M$ (-RA); or $M_{\varphi}=M^{2}$ patches along azimuth and $M_{r}=1$ patch along radius (-A). Note that we break this rule when the input resolution across radius or azimuth becomes lower than $M$ to maintain the number of patches in each window equals to $M^{2}$ at each stage (see tab. 1). We experimentally found (see sec. 5.6) that computing attention across $M^{2}$ patches along the azimuth (-A) yielded improved performance.

### v5

@@ Abstract
Wide-angle lenses are commonly used in perception tasks requiring a large field of view. Unfortunately, these lenses produce significant distortions making conventional models that ignore the distortion effects unable to adapt to wide-angle images. In this paper, we present a novel transformer-based model that automatically adapts to the distortion produced by wide-angle lenses. We leverage the physical characteristics of such lenses, which are analytically defined by the radial distortion profile (assumed to be known), to develop a distortion aware radial swin transformer (DarSwin). In contrast to conventional transformer-based architectures, DarSwin comprises a radial patch partitioning, a distortion-based sampling technique for creating token embeddings, and an angular position encoding for radial patch merging. We validate our method on classification tasks using synthetically distorted ImageNet data and show through extensive experiments that DarSwin can perform zero-shot adaptation to unseen distortions of different wide-angle lenses. Compared to other baselines, DarSwin achieves the best results (in terms of Top-1 accuracy) with significant gains when trained on bounded levels of distortions (very-low, low, medium, and high) and tested on all including out-of-distribution distortions. The code and models are publicly available at https://lvsn.github.io/darswin/
@@ 1 Introduction
Figure 1: Illustration of (cartesian) Swin [28] (left) and our (radial) DarSwin (right) given a wide-angle image (middle). While Swin [28] computes attention on the predefined windows over square image patches (bottom left orange region), DarSwin performs radial transformations using distortion-aware radial patches and computes the attention on windows defined over radial patches (shown in orange region bottom right), which enables greater generalization capabilities across different lenses.
@@
Unfortunately, such wide angle lenses create significant distortions in the image since the perspective projection model no longer applies: straight lines appear curved, and the appearance of the same object changes as a function of its position on the image plane. This distortion breaks the translation equivariance assumption implicit in convolution neural networks (CNNs) and therefore limits their applicability. This problem is further exacerbated by the diversity in lens distortion profiles: as we will demonstrate, a network naively trained for a specific lens tends to overfit to that specific distortion, and does not generalize well when tested on another lens. Just as methods are needed to address the “domain gap” [16] from dataset bias [42, 20], we believe we must also bridge the “distortion gap” to truly make wide angle imaging applicable.
@@
One popular strategy to bridge the distortion gap is to cancel the effect of distortion on the image plane by warping the input image back to a perspective projection model according to calibrated lens parameters. Conventional approaches can then be trained and tested on the resulting “un-distorted” images. A wide array of such methods, ranging from classical [5, 35, 52, 30, 19] to deep learning [47, 45], have been proposed. Unfortunately, warping a very wide angle image to a perspective projection tends to create severely stretched images and restricts the maximum field of view since, in the limit, a point at 90°90°90  ° azimuth projects at infinity. Reducing the maximum field of view defeats the purpose of using a wide angle lens in the first place. Other projections are also possible (e.g., cylindrical [33]) but these also tend to create unwanted distortions.
@@
Recently, methods that break free from the “undistort first” strategy aim to reason directly about the images without undistorting them. For example, methods like [34, 1] use deformable convolutions [8, 54] to adapt convolution kernels to the lens distortion. However, the high computational cost of deformable CNNs constrains the kernel adaptation to a few layers inside the network. Other approaches like spherical CNNs [6] or gauge equivariant CNN [7] can adapt to different manifolds but their applicability for lens distortion has not been demonstrated. Finally, vision transformers [11] and their more recent variants [53, 28, 44] could also better bridge the “distortion gap” since they do not assume any prior structure other than permutation equivariance, but their cartesian partitioning of the image plane do not take lens geometry into account (see fig. 1).
@@
In this paper, we present DarSwin, a transformer-based architecture that adapts its structure to the lens distortion profile, which is assumed to be known (i.e. the camera is calibrated). Our method, inspired by the recent Swin transformer architecture [28], leverages a distortion aware sampling scheme for creating token embeddings, employs polar patch partitioning and merging strategies, and relies on angular relative positional encoding. This explicitly embeds knowledge of the lens distortion in the architecture and makes it much more robust to the “distortion gap” created by training and testing on different lenses.
The main contribution of this paper is a novel transformer-based encoder that automatically adapts to the (known) lens distortion profile, which relies on a combination of the following novel distortion aware components: polar patch partitioning, distortion-based sampling scheme for creating token embeddings along with a jittering technique for better generalization, and angular relative position encoding for radial patch merging. We show, through extensive classification experiments, that DarSwin can perform zero-shot adaptation (without pretraining) across different lenses. Indeed, when our method is trained on a restricted set of distortions, we observe that it is much more robust to changes in distortions at test time than all of the compared alternatives, including baseline Swin [28] (applied on both distorted and undistorted images) and deformable attention transformer  [44].
@@ 2 Related work
Panoramic images span the full 360°360°360  ° field of view and are most commonly projected onto a plane using an equirectangular projection creates severe distortions especially near the poles. Many works have explicitly designed approaches to deal with equirectangular distortion, including depth estimation [55], saliency detection [50], segmentation [51], layout estimation [12], and object detection [39] to name a few. However, as with [6], these are specifically designed for spherical distortion and do not generalize to wide-angle lenses.
@@
Applying tasks like classification and object recognition on wide-angle images is relatively recent [36, 37, 33, 46, 49] due to the presence of distortions in the image. In computer vision, the application of wide-angle images ranges from visual perception [25] to autonomous vehicle cameras [49, 24, 26]. In this respect, the initial studies mainly focused on correcting the distortion of the image [45, 10, 47, 52, 21]. Recently, many convolutional-based and attention-based models have been proposed that try to directly reason on wide-angle images without relying on distortion correction models.
@@
CNNs [23, 40, 27] are particularly well-suited for perspective images due to their implicit bias and translational equivariance [4]. Methods like [41, 37, 36] tries to adapt CNNs on fisheye images for tasks such as object detection. However, the distortion caused by wide angle images breaks this symmetry, which reduces the generalization performance of CNNs. Deformable convolutions [8] (later extended in [54]) learn a deformation to be applied to convolution kernels which can provide greater flexibility at the cost of significant additional computation. Closer to our work, [1, 34, 9] use such deformable CNNs to understand the distortion in a fisheye image. In contrast, our DarSwin leverages attention-based mechanisms rather than convolution. Recently, Jang et al. [18] proposed a framework for distortion-aware domain adaptation, where a generator network is trained to transform an image to a different lens profile. In contrast, our method can adapt to a different lens without additional training.
@@
Vision transformers (ViT) [11] use self-attention mechanisms [43] computed on image patches rather than performing convolutions. Unlike CNNs, a ViT does not have a fixed geometric structure in its architecture: any extra structure is given via positional encoding. More recently, the Swin transformer architecture [28] proposes a multi-scale strategy of window-based attention. Later, deformable attention transformer (DAT) [44], adapts the concept of deformable CNNs [8, 54] to increase adaptability. Unlike existing transformer architectures, DarSwin explicitly embeds the distortion into its structure using polar sampling, patch partition, window-based self-attention, and angular positional encoding.
@@ 3 Image formation
Figure 2: Overview of our distortion aware transformer architecture, DarSwin (Azimuth Merge). It employs hierarchical layers of Swin-S transformer blocks [28] interspersed with patch merging layers. To make it adapt to lens distortion, the patch partition, linear embedding, and patch merging layers all take the lens projection curve $\mathcal{P}$ (c.f. sec. 3) as input.
@@ 4.1 Overview
Fig. 2 shows an overview of our proposed distortion aware transformer architecture. It accepts as input a single image and its distortion parameters in the form of a lens projection curve $\mathcal{P}(\theta)$ (see eq. 2). The image domain is first segmented into patches according to a polar partitioning module (sec. 4.2). Then, a linear embedding is computed from sampled points (sec. 4.3), reshaped into a radial-azimuth projection and fed to the first Swin transformer block. The attention mechanism employs an angular relative positional encoding scheme guided by the lens curve. This is followed by three blocks performing patch merging (sec. 4.6) and additional Swin transformer blocks. More details on each of these steps are provided below.
@@
Figure 3: Illustration of the difference between (a) Swin [28], which uses Cartesian image patches (shown in grey) and windows (shown in white borders) by making divisions along image axes, Swin performs attention on a Cartesian window and merges 2 $\times$ 2 neighbourhood patches to build a hierarchical structure (shown in the green and orange shaded region) (b) our proposed DarSwin uses polar image patches by making divisions along radius and azimuth. In our case, the images patches (shown in grey) and windows (shown in white borders) are defined along azimuth and merged along azimuth as (shown in the green and orange shaded region)
@@ 4.2 Polar partition
The first step of our proposed architecture is to partition the image domain (defined by a 2D plane) into patches. As opposed to Swin which performs the split in cartesian coordinates (fig. 3a), DarSwin employs a polar patch partitioning strategy (fig. 3b). After centering a polar coordinate system on the image center (assumed to be known), we first split according to azimuth angle $\varphi$ (in the image plane) in $N_{\varphi}$ equiangular regions. For the radial dimension, we split the image into $N_{r}$ radial regions such that the splits are equiangular in $\theta$ and obtain the corresponding radii using the lens projection function $\mathcal{P}(\theta)$ (see eqs. 2 and 4). The total number of patches is therefore $N_{\varphi}\times N_{r}$. In our experiments, we set $N_{r}=16$ and $N_{\varphi}=64$ for an input image of size $64\times 64$ pixels.
Figure 4: Example of distortion aware polar patch partition. Here, $N_{r}=4$ and $N_{\varphi}=4$ partitions along radius and azimuth respectively are used to illustrate (right, only the top-right quadrant of the image is shown). While the azimuth partitioning is performed in an equiangular fashion, the radial dimension takes the lens distortion curve (left) into account. The field of view along the incident angle $\theta$ is split into $N_{r}$ equal parts (left) and corresponding radial are obtained from the distortion curve. Hence for different lenses we can have different radii depending on the distortion parameters.
@@ 4.3 Linear embedding
The resulting image patches created by this approach have unequal amount of pixels. Therefore, we rely on a distortion aware sampling strategy to obtain the same number of points for each patch. To sample from the images, we define the number of sampling points along the radius and azimuth, as shown in fig. 5, and adapt the pattern according to each partition. In our experiments, we set 10 sampling points along radius and azimuth for each patch. Points are sampled in an equiangular fashion along the azimuth direction. For the radial dimension, we sample according to the same pattern as the polar partitioning (sec. 4.2); that is, we split in an equiangular fashion according to $\theta$ and obtain the corresponding radii using $\mathcal{P}(\theta)$ (eq. 2) as shown in fig. 5. The input image is sampled using bilinear interpolation, and samples are arranged in a polar format as illustrated in fig. 5. The resulting sample values are then fed into a linear embedding layer to produce token embeddings as input for the transformer block.
Figure 5: Example of sampling strategy on patch partitions with a total of 16 patches (here, the number of divisions along radius and azimuth are $N_{r}=2$ and $N_{\varphi}=8$ respectively) with window size $(1,2)$, i.e. ($M_{r}=1,M_{\varphi}=2$). We use bilinear interpolation to sample RGB values from the image (6 blue dots per patch) and arrange them in polar coordinates.
@@ 4.4 Window-based self attention
The Swin transformer architecture [28] uses window-based self-attention, where attention is computed on non-overlapping windows of $M\times M$ patches. Here, we maintain that strategy, but the polar nature of our patches allows for an additional design choice. Namely, the number of patches along azimuth $M_{\varphi}$ and radius $M_{r}$ can be different. As shown in tab. 1, we define two variants: $M$ patches along both azimuth and radius, $M_{r}=M_{\varphi}=M$ (DarSwin-RA); or $M_{\varphi}=M^{2}$ patches along azimuth and $M_{r}=1$ patch along radius (DarSwin-A). Note that we break this rule when the input resolution across radius or azimuth becomes lower than $M$ to maintain the number of patches in each window equals to $M^{2}$ at each Stage (see tab. 1). We experimentally found (see sec. 5.4) that computing attention across $M^{2}$ patches along the azimuth (DarSwin-A) yielded improved performance.